# The State of Open Science in Software Engineering Research: A Case Study of ICSE Artifacts

### v1

@@ Abstract
Replication packages are crucial for enabling transparency, validation, and reuse in software engineering (SE) research. While artifact sharing is now a standard practice and even expected at premier SE venues such as ICSE, the practical usability of these replication packages remains underexplored. In particular, there is a marked lack of studies that comprehensively examine the executability and reproducibility of replication packages in SE research. In this paper, we aim to fill this gap by evaluating 100 replication packages published as part of ICSE proceedings over the past decade (2015–2024). We assess the (1) executability of the replication packages, (2) efforts and modifications required to execute them, (3) challenges that prevent executability, and (4) reproducibility of the original findings. We spent approximately 650 person-hours in total executing the artifacts and reproducing the study findings. Our findings reveal that only 40% of the 100 evaluated artifacts were executable, of which 32.5% (13 out of 40) ran without any modification. Regarding effort levels, 17.5% (7 out of 40) required low effort, while 82.5% (33 out of 40) required moderate to high effort to execute successfully. We identified five common types of modifications and 13 challenges leading to execution failure, spanning environmental, documentation, and structural issues. Among the executable artifacts, only 35% (14 out of 40) reproduced the original results. These findings highlight a notable gap between artifact availability, executability, and reproducibility. Our study proposes three actionable guidelines to improve the preparation, documentation, and review of research artifacts, thereby strengthening the rigor and sustainability of open science practices in SE research.
I Introduction
@@
Open science promotes transparency, accessibility, and reusability by encouraging researchers to share publications, data, code, and methods with the community [Mendez2020OpenScience]. Backed by mandates from governments, funding organizations, and publishers [OECD2015, RoyalSociety2012], it has transformed fields like the natural and social sciences by enhancing reproducibility, reducing redundancy, and accelerating progress [Mirowski2018, Pontika2015]. Although software engineering (SE) research is highly empirical and tool-centric, making it especially well-suited for open science practices, the adoption of open science in SE has been slow and inconsistent [Mendez2019]. Nevertheless, in recent years, leading SE conferences and journals have taken steps to promote open science by introducing policies and artifact evaluation tracks, offering dedicated reviews and badges to incentivize the sharing of data, tools, and code [Ernst2021]. A recent study shows encouraging trends: 67.7% of SE papers from top-tier conferences (e.g., ICSE, FSE, ASE) between 2017 and 2022 included artifacts [Liu2024]. However, their actual usability in terms of executability and reproducibility, which are core aspects of open science, remains uncertain [Nong2023]. When curated effectively, these artifacts can enhance trust, enable reuse, and support rigorous validation and comparison across studies [Mendez2019]. Thus, a systematic investigation into the state of open science in SE is warranted to assess current practices, identify challenges to executability and reproducibility, and guide future improvements.
@@
Several studies explored open science in SE through artifact metadata inspection, surveys, literature reviews, and methodology analysis [Solari2018LabPackages, Cordeiro2025a, Mahmood2018, Gundersen2018Reproducibility]. For example, Solari et al. [Solari2018LabPackages] analyzed 26 lab packages to propose a reference model for artifact structure, highlighting common organizational patterns. Cordeiro et al. [Cordeiro2025a] surveyed SE researchers to understand reproducibility practices in controlled experiments. Rodríguez-Pérez et al. [RodriguezPerez2018Reproducibility] systematically reviewed 187 SZZ-related studies, which assessed reproducibility and reported limitations. In a broader context, Gundersen et al. [Gundersen2018Reproducibility] analyzed 400 AI papers to evaluate the availability of source code and the quality of documentation required for reproducibility. However, they did not attempt to execute or reproduce the studies. Going further, Nong et al. [Nong2023] executed 11 tools from 55 papers but focused narrowly on deep learning-based vulnerability detection in C/C++. While these efforts offer valuable insights, they primarily rely on secondary sources, with limited direct engagement in executing and evaluating artifacts. As a result, a comprehensive and hands-on assessment of open science practices across the broader SE research landscape remains an open challenge.
@@
In this study, we present the first large-scale empirical assessment of open science practices in SE by analyzing the executability, and reproducibility of 100 replication packages from International Conference on Software Engineering (ICSE) papers published over the past decade. First, we collected all ICSE research-track papers published from 2015 to 2024 and systematically inspected each paper to determine whether it provided a replication package link. For those who included a link, we visited it to see if it is still alive and accessible. We then reviewed the contents of each accessible package to determine whether it included executable components. Second, from the set of packages with executable components, we selected 100 replication packages using stratified random sampling [koyuncu2009ratio] and executed them in an isolated environment to systematically assess the degree of success and the effort required to achieve it. Third, for successfully executed packages, we compared the obtained results with those reported in the original studies to assess reproducibility. We spent approximately 650 person-hours executing the replication packages and analyzing their reproducibility. In particular, we made four major contributions by answering the following four research questions.
@@
Figure 1: Example executability issue in a replication package.
@@
Full Replication Package of our study is available in our online appendix [ReplicationPackage].
II Motivating Example
Despite growing efforts to promote open science, a notable gap exists between the availability of replication packages and their practical usability (e.g., executability and reproducibility). Consider two real-world examples that illustrate this gap. Figure 1 shows an open issue [seamfuzz_issue1] from an ICSE 2023 paper [Lee2023SEAMFUZZ], where a user attempted to run the shared artifact but encountered critical errors possibly caused by missing files and/or undocumented environmental assumptions. Notably, this artifact had earned both the available and reusable badges[acm_artifact_review_2024], suggesting that current artifact evaluation criteria may not fully reflect the practical usability challenges. Figure 2 presents another case [ltlfuzzer_issue8] from ICSE artifact [Meng2022LTTFuzz], where a user reported that the code was not usable out-of-the-box due to missing required updates that had not been pushed to the public repository. As a result, the replication attempt failed despite the artifact being publicly available and endorsed.
Figure 2: Example of a user’s concern about the quality of the artifact.
@@ II Motivating Example
TABLE I: Summary of ICSE Papers, Artifact Availability, and Executability (2015–2024)
@@
III Study Methodology
III-A Dataset Construction
@@
Corpus Collection. We collected 1,372 ICSE research track papers published from 2015 to 2024 from publication databases (e.g., IEEE, ACM) using their Digital Object Identifiers. Table I shows the yearly distribution of ICSE papers with replication packages, accessibility, and executability readiness.
@@
Artifact Identification Process. To determine whether a paper shared replication package links, we employed the following three-pass search strategy. In Pass 1, we searched for open science–related keywords (e.g., replication, reproduce, artifact, package, dataset) in the PDF to identify artifact links. To construct this keyword list, we manually examined an initial set of 20 papers, which resulted in a list of 25 terms (please refer to our replication package [ReplicationPackage]) used consistently throughout the analysis. If no link was identified via keyword search, in Pass 2, we manually scanned the full text, focusing on the end of the introduction, dedicated sections (e.g., Data Availability), footnotes, and the reference list to locate any implicit mentions of artifacts or URLs. For papers that discussed replication packages but did not include explicit artifact links, in Pass 3, we conducted a web search using the paper title and author names to check for artifacts hosted on platforms such as GitHub, Zenodo, or institutional websites. Interestingly, we found one artifact link in this process. Among 1,372 ICSE papers, 1,085 (79.08%) included at least one artifact link. While some artifacts were hosted on multiple platforms (21 cases), we collected all provided links to assess their availability.
@@
Artifact Availability with Executable Components. For papers that included artifact links, we applied two filtering criteria: (1) availability – the provided links were active and the artifact contents could be accessed at the time of inspection, and (2) executability – the artifacts contained executable components (e.g., code, scripts, or tools) to initiate an execution attempt. Based on these filtering criteria, we identified 941 papers with available artifact links and 796 papers whose artifacts were both available and included executable components. Among 1372 papers, 287 did not provide any replication package link, and 144 provided links that were inaccessible.
@@
Sampling Strategy. To ensure temporal representativeness, we applied year-wise stratified random sampling [koyuncu2009ratio] to select 100 artifacts for execution and reproducibility analysis, as summarized in Table I. The choice of 100 packages balanced breadth and feasibility, ensuring wide coverage of SE studies while keeping the manual execution and assessment workload practical yet rigorous. We initially planned to sample approximately 10% from each year (around 80 packages), but expanded this to 100 to enhance diversity and strengthen the validity of our findings. Although not based on a formal power analysis, this sample size is consistent with standard practice in empirical SE research, where extensive manual evaluation typically limits the proportion of artifacts analyzed while still supporting representative and reliable conclusions.
III-B Assessing Executability, Effort, and Required Modifications
@@
III-B1 Retrieving and Preparing Artifacts
@@
III-B2 Setting up Isolated Execution Environments
@@
$\bullet$ Handling Missing Specifications. Our primary source for system specifications was the artifact’s documentation. However, when artifacts did not specify system requirements, we used our M2 machine (Table II) as the default hardware configuration. To approximate the OS version, we applied a two-year backward buffer from the artifact’s publication year, considering the typical gap between the start of the experiment and the publication date. For example, for an artifact published in 2020, we used the latest long-term support OS version available in 2018.
III-B3 Installing Dependencies and External Tools
@@ III-B4 Running the Experiment
TABLE II: Machine Details
@@
III-B5 Troubleshooting and Fixing Runtime Errors
@@
Once modifications were applied, we returned to the previous execution step (Section III-B4) and re-ran the experiment. If the issues were resolved, we continued with the remaining execution steps. Otherwise, we resumed troubleshooting. Troubleshooting was often time-consuming, with the number and complexity of errors varying across artifacts. To manage effort effectively, we set a maximum effort threshold of four hours. If an issue could not be resolved within that window and the artifact lacked sufficient documentation, we stopped further investigation and proceeded to the final execution decision.
III-B6 Labeling
@@
III-C Analyzing Failure Causes and Executability Challenges
@@
III-D Examining Reproducibility of Executable Artifacts and the Factors Behind Failure
@@
III-E Experimental Setup
To conduct the execution and reproducibility experiments, the first two authors analyzed artifacts using a total of seven machines, selected to closely align with the requirements specified in the replication packages. Table II provides an overview of the hardware and operating system specifications used in this study. Each evaluator had access to at least one Windows machine for running artifacts that provided Windows-specific instructions or executables, and two Ubuntu-based Linux machines, at least one of which was equipped with a GPU to support experiments involving hardware acceleration or deep learning workloads. We kept a Mac machine as a backup and needed it only for one artifact that required Mac-specific data processing.
IV Study Findings
@@
IV-A Answering RQ1: Executability Status, Level of Effort and Modifications to Execute Replication Package
[…]
Figure 3: Executability status of artifacts over the years.
@@ IV-A Answering RQ1: Executability Status, Level of Effort and Modifications to Execute Replication Package
Level of Effort Required to Execute Replication Packages. Figure 4(a) presents the relationship between executability status and the level of effort required to execute the artifacts. The results show a stark contrast in effort distribution across executable, partially executable, and non-executable artifacts. Among the 40 executable artifacts, the majority (72.5%) required either low (11) or moderate (18) effort. However, 11 executable artifacts (27.5%) still demanded high effort, indicating that even runnable artifacts often lack sufficient documentation, packaging, or automation support. The situation is more severe for partially executable artifacts (24). All required moderate (13) or high (11) effort, yet none could be executed successfully. Even for the non-executable artifacts (36), we invested non-trivial time: 52.78% (19/36) required moderate to high effort before reaching a dead end. The remaining 47.22% (17/36) failed early due to severe deficiencies or incompleteness and were classified under low effort, not because they were simple, but because they were unworkable from the start. These results emphasize the urgent need for community-wide practices that minimize unnecessary overhead by promoting thorough environment specification, proper automation, and quality assurance of replication packages.
@@
Figure 4: Effort & Modification Vs. Executability
IV-B Answering RQ2: Types of Modifications
We conducted a qualitative analysis (as discussed in Section III-B5) and labeled the observed modifications using a hierarchical classification spanning five broad categories and 14 specific types.

### v2

@@ Abstract.
Replication packages are crucial for enabling transparency, validation, and reuse in software engineering (SE) research. While artifact sharing is now a standard practice and even expected at premier SE venues such as ICSE, the practical usability of these replication packages remains underexplored. In particular, there is a marked lack of studies that comprehensively examine the executability and reproducibility of replication packages in SE research. In this paper, we aim to fill this gap by evaluating 100 replication packages published in ICSE proceedings over the past decade (2015–2024). We assess the (1) executability of the replication packages, (2) efforts and modifications required to execute them, (3) challenges that prevent executability, and (4) reproducibility of the original findings. We spent approximately 650 person-hours in total executing the artifacts and reproducing the study findings. Our findings reveal that only 40% of the 100 evaluated artifacts were executable, of which 32.5% (13 out of 40) ran without any modification. Regarding effort levels, 17.5% (7 out of 40) required low effort, whereas 82.5% (33 out of 40) required moderate to high effort to be executed successfully. We identified five common types of modifications and 13 challenges that lead to execution failure, encompassing environmental, documentation, and structural issues. Among the executable artifacts, only 35% (14 out of 40) reproduced the original results. These findings highlight a notable gap between artifact availability, executability, and reproducibility. Our study proposes three actionable guidelines to improve the preparation, documentation, and review of research artifacts, thereby strengthening the rigor and sustainability of open science practices in SE research.
Open Science, Executability, Reproducibility, Replication Package, Software Engineering
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazilisbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Empirical software engineeringccs: Software and its engineering Reusabilityccs: General and reference Evaluation
1. Introduction
@@
Open science promotes transparency, accessibility, and reusability by encouraging researchers to share publications, data, code, and methods with the community (Mendez et al., 2020). Backed by mandates from governments, funding organizations, and publishers (OECD, 2015; The Royal Society, 2012), it has transformed fields like the natural and social sciences by enhancing reproducibility, reducing redundancy, and accelerating progress (Mirowski, 2018; Pontika et al., 2015). Although software engineering (SE) research is highly empirical and tool-centric, making it especially well-suited for open science practices, the adoption of open science in SE has been slow and inconsistent (Méndez Fernández et al., 2019). Nevertheless, in recent years, leading SE conferences and journals have taken steps to promote open science by introducing policies and artifact evaluation tracks, offering dedicated reviews and badges to incentivize the sharing of data, tools, and code (Ernst et al., 2021). A recent study shows encouraging trends: 67.7% of SE papers from top-tier conferences (e.g., ICSE, FSE, ASE) between 2017 and 2022 included artifacts (Liu et al., 2024). However, their actual usability in terms of executability and reproducibility, which are core aspects of open science, remains uncertain (Nong et al., 2023). When curated effectively, these artifacts can enhance trust, enable reuse, and support rigorous validation and comparison across studies (Méndez Fernández et al., 2019). Thus, a systematic investigation into the state of open science in SE is warranted to assess current practices, identify challenges to executability and reproducibility, and guide future improvements.
@@
Several studies have explored open science in SE through the inspection of artifact metadata, surveys, literature reviews, and methodological analyses (Solari et al., 2018; Cordeiro and Oliveira Jr, 2025a; Mahmood et al., 2018; Gundersen and Kjensmo, 2018b). For example, Solari et al. (Solari et al., 2018) analyzed 26 lab packages to propose a reference model for artifact structure, highlighting common organizational patterns. Cordeiro et al. (Cordeiro and Oliveira Jr, 2025a) surveyed SE researchers to understand reproducibility practices in controlled experiments. Rodríguez-Pérez et al. (Rodríguez-Pérez et al., 2018b) systematically reviewed 187 SZZ-related studies, which assessed reproducibility and reported limitations. In a broader context, Gundersen et al. (Gundersen and Kjensmo, 2018b) analyzed 400 AI papers to evaluate the availability of source code and the quality of documentation required for reproducibility. However, they did not attempt to execute or reproduce the studies. Going further, Nong et al. (Nong et al., 2023) executed 11 tools from 55 papers but focused narrowly on deep learning-based vulnerability detection in C/C++. While these efforts offer valuable insights, they primarily rely on secondary sources, with limited direct engagement in executing and evaluating artifacts. As a result, a comprehensive and hands-on assessment of open science practices across the broader SE research landscape remains an open challenge.
@@
In this study, we present the first large-scale empirical assessment of open science practices in SE by analyzing the executability, and reproducibility of 100 replication packages from International Conference on Software Engineering (ICSE) papers published over the past decade. First, we collected all ICSE research-track papers published from 2015 to 2024 and systematically inspected each paper to determine whether it provided a replication package link. For those who included a link, we visited it to see if it is still alive and accessible. We then reviewed the contents of each accessible package to determine whether it included executable components. Second, from the set of packages with executable components, we selected 100 replication packages using stratified random sampling (Koyuncu and Kadilar, 2009) and executed them in an isolated environment to systematically assess the degree of success and the effort required to achieve it. Third, for successfully executed packages, we compared the obtained results with those reported in the original studies to assess reproducibility. We spent approximately 650 person-hours executing the replication packages and analyzing their reproducibility. In particular, we made four major contributions by answering the following four research questions.
@@
Full Replication Package of our study is available in our online appendix (Replication Package, 2026).
Figure 1. Example of an issue in a replication package, where users encountered critical execution errors.
A screenshot of a GitHub issue showing error messages and stack traces encountered when attempting to execute a replication package.
2. Motivating Example
Despite growing efforts to promote open science, a notable gap exists between the availability of replication packages and their practical usability (e.g., executability and reproducibility). Consider two real-world examples that illustrate this gap. Figure 1 illustrates an executability issue in a replication package. In particular, it highlights an open issue (SeamFuzz Contributors, 2024) from an ICSE 2023 paper (Lee et al., 2023), in which a user attempted to run the shared artifact but encountered critical errors that may have been caused by missing files and/or undocumented environmental assumptions. Notably, this artifact earned both the available and reusable badges (Association for Computing Machinery, 2024), suggesting that current artifact evaluation criteria may not fully reflect practical usability challenges. Figure 2 presents another case (LTL-Fuzzer Contributors, 2024) from the ICSE artifact (Meng et al., 2022), where a user reported that the code was not usable out-of-the-box due to missing required updates that had not been pushed to the public repository. As a result, the replication attempt failed despite the artifact being publicly available and endorsed.
Figure 2. Example of a user’s concern about the quality of the artifact.
A screenshot showing user concerns about artifact quality, including comments about incomplete documentation, missing repository updates, and difficulty understanding how to execute the replication package.
@@ 2. Motivating Example
Table 1. Summary of ICSE Papers, Artifact Availability, and Executability (2015–2024)
@@
3. Study Methodology
3.1. Dataset Construction
@@
Corpus Collection. We collected 1,372 ICSE research track papers published from 2015 to 2024 from publication databases (e.g., IEEE, ACM) using their Digital Object Identifiers. Table 1 shows the yearly distribution of ICSE papers with replication packages, accessibility, and executability readiness.
@@
Artifact Identification Process. To determine whether a paper shared replication package links, we employed the following three-pass search strategy. In Pass 1, we searched for open science–related keywords (e.g., replication, reproduce, artifact, package, dataset) in the PDF to identify artifact links. To construct this keyword list, we manually examined an initial set of 20 papers, which resulted in a list of 25 terms (please refer to our replication package (Replication Package, 2026)) used consistently throughout the analysis. If no link was identified via keyword search, in Pass 2, we manually scanned the full text, focusing on the end of the introduction, dedicated sections (e.g., Data Availability), footnotes, and the reference list to locate any implicit mentions of artifacts or URLs. For papers that discussed replication packages but did not include explicit artifact links, in Pass 3, we conducted a web search using the paper title and author names to check for artifacts hosted on platforms such as GitHub, Zenodo, or institutional websites. Interestingly, we found one artifact link in this process. Among 1,372 ICSE papers, 1,085 (79.08%) included at least one artifact link. While some artifacts were hosted on multiple platforms (21 cases), we collected all provided links to assess their availability.
@@
Artifact Availability with Executable Components. For papers that included artifact links, we applied two filtering criteria: (1) availability – the provided links were active, and the artifact contents could be accessed at the time of inspection, and (2) executability – the artifacts contained executable components (e.g., code, scripts, or tools) to initiate an execution attempt. Based on these filtering criteria, we identified 941 papers with available artifact links and 796 papers whose artifacts were both available and included executable components. Among 1372 papers, 287 did not provide any replication package link, and 144 provided links that were inaccessible.
@@
Sampling Strategy. To ensure temporal representativeness, we applied year-wise stratified random sampling (Koyuncu and Kadilar, 2009) to select 100 artifacts for execution and reproducibility analysis, as summarized in Table 1. The choice of 100 packages balanced breadth and feasibility, ensuring wide coverage of SE studies while keeping the manual execution and assessment workload practical yet rigorous. We initially planned to sample approximately 10% from each year (around 80 packages), but expanded this to 100 to enhance diversity and strengthen the validity of our findings. Although not based on a formal power analysis, this sample size is consistent with standard practice in empirical SE research, where extensive manual evaluation typically limits the proportion of artifacts analyzed while still supporting representative and reliable conclusions.
3.2. Assessing Executability, Effort, and Required Modifications
@@
3.2.1. Retrieving and Preparing Artifacts
@@
3.2.2. Setting up Isolated Execution Environments
@@
$\bullet$ Handling Missing Specifications. Our primary source for system specifications was the artifact’s documentation. However, when artifacts did not specify system requirements, we used our M2 machine (Table 2) as the default hardware configuration. To approximate the OS version, we applied a two-year backward buffer from the artifact’s publication year, considering the typical gap between the start of the experiment and the publication date. For example, for an artifact published in 2020, we used the latest long-term support OS version available in 2018.
3.2.3. Installing Dependencies and External Tools
@@ 3.2.4. Running the Experiment
Table 2. Machine Details
@@
3.2.5. Troubleshooting and Fixing Runtime Errors
@@
Once modifications were applied, we returned to the previous execution step (Section 3.2.4) and re-ran the experiment. If the issues were resolved, we continued with the remaining execution steps. Otherwise, we resumed troubleshooting. Troubleshooting was often time-consuming, with the number and complexity of errors varying across artifacts. To manage effort effectively, we set a maximum effort threshold of four hours. If an issue could not be resolved within that window and the artifact lacked sufficient documentation, we stopped further investigation and proceeded to the final execution decision.
3.2.6. Labeling
@@
3.3. Analyzing Failure Causes and Executability Challenges
@@
3.4. Examining Reproducibility of Executable Artifacts and the Factors Behind Failure
@@
3.5. Experimental Setup
To conduct the execution and reproducibility experiments, the first two authors analyzed artifacts using a total of seven machines, selected to closely align with the requirements specified in the replication packages. Table 2 provides an overview of the hardware and operating system specifications used in this study. Each evaluator had access to at least one Windows machine for running artifacts that provided Windows-specific instructions or executables, and two Ubuntu-based Linux machines, at least one of which was equipped with a GPU to support experiments involving hardware acceleration or deep learning workloads.
We kept a Mac machine as a backup and needed it only for one artifact that required Mac-specific data processing.
4. Study Findings
@@
4.1. Answering RQ1: Executability Status, Level of Effort and Modifications to Execute Replication Package
Figure 3. Executability status of artifacts over the years.
A stacked bar chart showing executability status from 2015 to 2024. Each bar represents one year and is divided into three segments: executable (green), partially executable (yellow), and not executable (red).
@@ 4.1. Answering RQ1: Executability Status, Level of Effort and Modifications to Execute Replication Package
Figure 4. Effort & Modification Vs. Executability
Two side-by-side stacked bar charts. The left chart shows effort levels (High, Medium, Low) required for execution. The right chart shows modification levels (Significant, Moderate, Minor, None) needed to make artifacts executable.
@@
Level of Effort Required to Execute Replication Packages. Figure 4(a) presents the relationship between executability status and the level of effort required to execute the artifacts. The results show a stark contrast in effort distribution across executable, partially executable, and non-executable artifacts. Among the 40 executable artifacts, the majority (72.5%) required either low (11) or moderate (18) effort. However, 11 executable artifacts (27.5%) still demanded high effort, indicating that even runnable artifacts often lack sufficient documentation, packaging, or automation support. The situation is more difficult for partially executable artifacts (24). All required moderate (13) or high (11) effort, yet none could be executed successfully. Even for the non-executable artifacts (36), we invested non-trivial time: 52.78% (19/36) required moderate to high effort before reaching a dead end. The remaining 47.22% (17/36) failed early due to severe deficiencies or incompleteness and were classified under low effort, not because they were simple, but because they were unworkable from the start. These results emphasize the urgent need for community-wide practices that minimize unnecessary overhead by promoting thorough environment specification, proper automation, and quality assurance of replication packages.
Figure 5. Types of Modifications
A hierarchical taxonomy chart showing types of modifications required for artifact execution. The chart has 5 main categories at the top level: Configuration, Dependency, Code, Environment, and Documentation.
@@
4.2. Answering RQ2: Types of Modifications Needed to Execute Replication Packages
We conducted a qualitative analysis (as discussed in Section 3.2.5) and labeled the observed modifications using a hierarchical classification spanning five broad categories and 14 specific types.